OpenStudy (anonymous):

Please help I'm so freakin confused!!!!! Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses.

OpenStudy (anonymous):

\[f(x)=\frac{ x+a }{ b} \] and \[g(x)=cx-d\]

OpenStudy (anonymous):

@campbell_st @hartnn I have absolutely no clue what this means!!

OpenStudy (anonymous):

@Destinymasha @radar